# Fast Low Energy Reconstruction using Convolutional Neural Networks

### v1

@@ 1 Introduction
Figure 1: The upper panel shows a top view of the IceCube detector. DeepCore strings are indicated by red circles and IceCube strings by green circles. The lower panel shows a side view of the detector. The green shading highlights the DeepCore DOMs and the red shading highlights the veto DOMs.
When neutrinos interact in or near the IceCube detector and produce charged particles, the relativistic charged particles propagate in the ice and give rise to Cherenkov photons. DOMs record data when the PMT signal voltage passes a threshold equivalent to one quarter the amplitude of an average photoelectron (PE) response [3]. To filter our random noise, a local-coincidence trigger is applied, requiring that signals from multiple DOMs occur within a predefined time window [4]. Once a trigger is formed, the PMT signals are processed by the analog to digital converters into digitized waveforms. Waveforms are converted into pulses using a decomposition into a basis of the known response of the PMT to a photoelectron. Using the charge and time information extracted from the pulses of each DOM, algorithms [5, 6, 7] are employed to reconstruct the neutrino interactions in the detector.
Convolutional neural networks (CNN) are deep learning models that excel in processing images. By learning features through convolutional layers, CNNs can efficiently extract patterns from complex data. In reconstruction applications for particle interactions in large detectors [8, 9, 5], CNNs are increasingly employed to interpret high-dimensional data and enhance accuracy in tasks such as event classification and signal processing. In previous IceCube reconstruction studies for high-energy astrophysical neutrinos, a CNN was developed to process data from all 86 strings [5]. In this work, we introduce a different approach that uses CNNs specifically optimized to reconstruct sub-100 GeV neutrino interactions concentrated in the DeepCore detector. The results of this low-energy CNN was employed in a recent atmospheric neutrino oscillation measurement [1], and the aim of this paper is to provide the further technical details that complement the published results.
@@ 2 Input data and Neural Network Architecture
The CNNs in this study are employed for the five distinct use cases to reconstruct low-energy neutrino interactions: the estimation of the neutrino’s energy, arrival direction, and interaction vertex position, and the binary classification between track-like signal events, $\nu_{\mu}$ charged-current (CC) interactions versus cascade-like background events, including $\nu_{\mu}$ neutral-current (NC) interactions, and background atmospheric muon events versus signal neutrino events. For GeV-scale neutrino reconstruction in DeepCore, it is most productive to use data from in and immediately around the DeepCore region, rather than the full detector. The CNNs use all eight DeepCore strings and the 19 surrounding IceCube strings, which form the two rings outside the DeepCore region, as shown in the orange hexagon in Figure 1. These two rings help optimize reconstruction performance in the DeepCore region by providing additional information from adjacent IceCube DOMs, while avoiding the computational cost of including farther DOMs, which typically do not provide meaningful data for reconstructing neutrino interactions within DeepCore. Due to differences in spatial distance and quantum efficiency between DOMs on the DeepCore and IceCube strings, data from these two categories are fed into separate input layers, each with a dimension of [Nstring, 60, 5], where Nstring represents the number of strings from which data is included (19 and 8 for the IceCube and DeepCore input layers respectively) and 60 corresponds to the 60 DOMs on each string. When the training datasets are processed, these are treated as Nstring × 60 images with 5 summarized variables (channels) calculated. The CNN kernel spans only the vertical, or $z$-depth, of the strings with the kernel size given in Figure 2, ranging from taking a maximum of 2 DOMs above and below the target DOM in the convolution. Since the DeepCore strings are deployed in an irregular array, no convolution is applied in the $xy$-plane. The same convolutional kernel applied to the IceCube strings to be consistent, and thus both branches of the CNN use the same kernel applied in $z$-depth only. After passing through functional layers, such as “Conv 2D", “Batch Normalization", “Max Pooling", and “Dropout" in the sub-networks (see Figure 2 for details), the outputs of the two sub-networks are concatenated and fed into one fully connected dense layer. This last layer is then used to estimate the physical properties, which serve as the labels for the CNN.
@@
The data from each DOM are summarized from the charge and time information of the digitized pulses recorded in the time window of the event interaction, typically within -500 to 4000 ns of the DeepCore trigger time [10]. This time window corresponds to about eight times the scattering length divided by the running timelight in the ice, and thus the pulses falling beyond this window are considered as noise. To ensure consistency across all data and simulations, we quantize the charges in steps of 0.05 PE before using them to calculate the CNN input features. Five summary variables are calculated from the pulse series for each DOM to represent the statistical features of the digitized waveform: total charge, time of the first pulse, time of the last pulse, charge weighted mean of pulse times, and charge weighted standard deviation of pulse times.
These summary variables have significantly different characteristic scales. To train the networks more efficiently, we rescale the input variables by dividing the expected maximum values of the input, which are listed as the division factors in Table 1.
@@
The same CNN architecture was trained for five distinct use cases, separately reconstructing different physical properties of a particle interaction in the detector. In the output layer, the CNNs use linear activation functions for estimating arrival direction (zenith angle) and interaction vertex coordinates ($x$, $y$, $z$), and rectified linear units (ReLU) for predicting neutrino energy. The difference is because zenith angle, in units of radians, and vertex coordinates, in units of meters, do not have hard boundaries while energy must be positive. For the binary classifications, a binary entropy loss function was used. Specialized training samples were used to achieve various training goals and ensure a balanced distribution of events in the training sets, as discussed in Section 4.
@@ 3 Loss function for CNN training
Table 2: Loss functions used for each of the networks, equations show in the loss function plots
@@
where $n$ is the number of events, $t_{i}$ is the true value and $r_{i}$ is the reconstructed (or estimated) variable by the CNN. This was chosen carefully such that the energy network would give equal importance to reconstructing the low energy events (at a few GeV) as the high energy events (at 100 GeV). Since the $\mathcal{O}$(10) GeV events are the region where oscillation is expected, the percentage error ensures that the 10 GeV events are well-reconstructed without over-emphasis on the $\mathcal{O}$(100) GeV events. The vertex and zenith networks use the mean squared error loss:
@@
The CNNs for classification purposes use binary cross-entropy loss function:
$\textrm{Loss}=-{(t\ln(r)+(1-t)\ln(1-r))},$ (3.3)
where $t$ is the MC truth of being signal (1) or background (0), and $r$ is the predicted probability of being a signal event.
@@ 4 Training Samples
Simulated Monte-Carlo (MC) datasets of neutrino interactions and atmospheric muons were used to train the CNNs. In the pre-processing, data quality cuts and random shuffling were applied to ensure the desired distributions were used to train for different CNNs, as discussed further below.
The neutrino MC datasets were generated with the GENIE package (version 2.12.8) [11]. This dataset has larger statistics and is independent from the MC events that were used in the neutrino oscillation data analyses [1]. The muon MC samples were generated with MuonGun, an IceCube software package that simulates atmospheric muons from cosmic rays (method based on [12]). The simulated muon events used for training were from a subset of what is used in the oscillation analysis. IceCube neutrino analyses typically involve filters, applied successively, to remove background-like events and keep high-quality neutrino candidates. The training sets for the CNNs were processed through the same sets of filters and basic reconstruction routines commonly used for DeepCore oscillation data analysis (see Section III in [4] for details). However, a boosted decision tree (BDT) [13] was trained using lower-level variables and applied to remove neutrino-like background events, mostly atmospheric muons, therefore noted as BDTμ (see Section III.C of  [4] for details). To drop the noise-like events and keep the track-like $\nu_{\mu}$ charged-current (CC) events, the score of the BDTμ was required to be at least 0.95 in all of the CNN training datasets, except for energy, zenith, and the atmospheric muon dataset used for training the atmospheric muon classification. To retain enough statistics, the atmospheric muon events used for training were taken from the nominal dataset before the BDTμ filter was applied.
@@
Energy $\nu_{\mu}$ CC 9 million 1-500
Zenith $\nu_{\mu}$ CC 5 million 1-300
Vertex $\nu_{\mu}$ CC 5 million 1-500
@@ 4.2 Neutrino Zenith Angle
A dataset of $\nu_{\mu}$ CC events between 5-300 GeV assuming an injection spectrum that is uniformly distributed in zenith angle in [0, $\pi$] was simulated for the zenith training. To produce a training dataset with a uniform distribution of zenith angles, events were rejected until each 0.01*$\pi$ bin contained a similar number of events. After processing, 5 million simulated events remained for training and testing. Training the zenith network, however, relies on showing the network the full track signature. Thus, a cut was applied so that the CNN training set contained only events where the paths of all of the secondary particles, produced in the neutrino interaction, were contained within the active detector volume region indicated by the orange hexagon in Figure 1. This cut is referred to as “containment" in Table 4 and corresponds to the variable restrictions ($-505$ m < $z$ < 500 m and within 260 m to the center-most string, $r_{\rm string36}$).
@@ 5 Performance
The performance is evaluated on the MC dataset used in the oscillation analysis described in [1] and listed in Table II in [4], with atmospheric flux, neutrino interactions, and oscillation weights applied. The previous IceCube neutrino oscillation result [4] used the SANTA tool for directional reconstruction and the LEERA tool for energy reconstruction, both described in Section III.B of [14]. We compared the performance of our CNN routines with these tools using a common subset of events that passed the selections of both the current and previous oscillation analyses. To show the performance on neutrino energy and zenith angle (see Figures 3 and  4), the cuts on energy or zenith angle are not applied. The SANTA algorithm requires events to pass the selection criteria described in Section 2 of [7], it is only applicable to a small set of high-quality events in the DeepCore detector. A more inclusive, likelihood-based algorithm called RETRO has been developed to provide more comprehensive information, allowing it to be applied to a larger set of events and yielding better performance [7]. However, the performance improvements achieved with RETRO come at the cost of significantly slower processing times compared to SANTA/LEERA. The CNNs presented in this paper, which are designed and optimized for the DeepCore region, address the slow processing time issue but offer comparable performance. A comparison of the runtime for SANTA, LEERA, RETRO and the CNNs discussed in this paper, is given in Table 5.recotimes A detailed comparison between CNN and RETRO reconstructions can be found in Appendix A.
@@ 5.1 Comparison to Reconstruction used in the Prior DeepCore Oscillation Result
Besides the improvement in reconstruction performance in the selected analysis samples, the CNN technique also shows an advantage in processing time (see Table. 5) compared to SANTA as well as RETRO. The graphics processing unit (GPU) used for evaluation was a Tesla V100S with a central processing unit (CPU) of Intel(R) Xeon(R) Platinum 8260 CPU @ 2.40 GHz. The same CPU was used to evaluate the inference speed the 5 CNNs. The precision of the CNN parameters is float32.
Table 5: Runtime of applying the different reconstructions to the MC sample where the last two columns are scaled from the first one to get a sensible estimate on a large statistics of production.
@@
5 CNNs on GPU 0.0011 8×$\times 10^{7}$ 2 min
5 CNNs on CPU 0.29 3×$\times 10^{5}$ 8 hours
SANTA/LEERA 0.16 [7] $\times 10^{5}$ 4 hours
RETRO 40 [7] 2×$\times 10^{3}$ 46 days
@@ 6 Summary
Ongoing efforts include adapting low-energy CNNs for reconstructing low-energy events, with a focus on azimuth and estimating reconstruction uncertainty. These developments will be valuable for future IceCube analyses in both particle physics and astrophysics. The IceCube collaboration continues to prioritize improving reconstruction resolution and runtime through the development of different algorithms. At TeV energies, a CNN takes the whole detector for cascade reconstruction into account [5]. Focusing on sub-100 GeV energies, we developed a CNN that uses DeepCore and surrounding IceCube strings. Additionally, efforts are underway to develop an atmospheric neutrino sample using graph neural networks [6].
@@ Appendix A Comparison to RETRO Reconstruction
Figure 9 shows the comparisons of CNN and RETRO on a subset of overlapping events that pass all the analysis cuts of the CNN-based sample and the likelihood fitting quality cut of RETRO with flux, neutrino interaction cross-section, and oscillation weights applied.
@@
The CNN reconstructs the interaction position in ($x$, $y$, $z$) coordinates, but the analysis cut is applied on depth ($z$) and horizontal radius relative to the position of the IceCube string-36 ($\rho_{36}$). Therefore, we show the performance of $z$, $\rho_{36}$, $\cos(\theta_{\rm{zenith}})$, and energy along the true neutrino energy as in Figure 9. In the $\nu_{\mu}$ disappearance analysis [1], the signal events are $\nu_{\mu}$ CC. To optimize reconstruction performance for these signal events, the CNNs for neutrino energy, zenith angle, and interaction vertex were trained and tested on $\nu_{\mu}$ CC events. Similar evaluation plots for the CNN reconstruction of cascade-like events, using $\nu_{e}$ CC as an example, can be found in Figure 10, which also shows reasonable performance.

### v2

@@ 1 Introduction
Figure 1: The upper panel shows a top view of the IceCube detector. DeepCore strings are indicated by red circles and IceCube strings by green circles. The DeepCore detector consists of both DeepCore strings and IceCube strings located within the dashed green hexagonal circle. In this work, the CNNs use only the strings within the orange hexagon. The lower panel shows a side view of the detector. The green shading highlights the DeepCore DOMs and the red shading highlights the veto DOMs.
When neutrinos interact in or near the IceCube detector and produce charged particles, the relativistic charged particles propagate in the ice and give rise to Cherenkov photons. DOMs record data when the PMT signal voltage passes a threshold equivalent to one quarter the amplitude of an average photoelectron (PE) response [12]. To filter out random noise, a local-coincidence trigger is applied, requiring that signals from multiple DOMs occur within a predefined time window [4]. Once a trigger is formed, the PMT signals are processed by the analog to digital converters into digitized waveforms. Waveforms are converted into pulses using a decomposition into a basis of templates of the known response of the PMT to a photoelectron. Using the charge and time information extracted from the pulses of each DOM, algorithms [5, 6, 13] are employed to reconstruct the neutrino interactions in the detector.
Convolutional neural networks (CNN) are deep learning models that excel in processing images. By learning features through convolutional layers, CNNs can efficiently extract patterns from complex data. In reconstruction applications for particle interactions in large detectors [9, 10, 5], CNNs are increasingly employed to interpret high-dimensional data and enhance accuracy in tasks such as event classification and signal processing. In previous IceCube reconstruction studies for high-energy astrophysical neutrinos, a CNN was developed to process data from all 86 strings [5]. In this work, we introduce a different approach that uses CNNs specifically optimized to reconstruct sub-100 GeV neutrino interactions concentrated in the DeepCore detector. The results of this low-energy CNN were employed in a recent atmospheric neutrino oscillation measurement [3], and the aim of this paper is to provide the further technical details that complement the published results.
@@ 2 Input data and Neural Network Architecture
The CNNs in this study are employed for the five distinct use cases to reconstruct low-energy neutrino interactions: the estimation of the neutrino’s energy, arrival direction, and interaction vertex position, and the binary classification between track-like signal events, $\nu_{\mu}$ charged-current (CC) interactions versus cascade-like background events, including $\nu_{\mu}$ neutral-current (NC) interactions, and background atmospheric muon events versus signal neutrino events. The same CNN architecture, with different output layers for different training goals, was trained for five distinct use cases, separately reconstructing different physical properties of a particle interaction in the detector. Another advantage of training five CNNs instead of a combined one is to optimize the performance of each training by employing specialized training samples to ensure a balanced distribution of events in the training sets, as discussed in Section 4. The shared architecture and the necessary difference in the output layers are discussed in detail in this section. For GeV-scale neutrino reconstruction in DeepCore, it is most productive to use data from in and immediately around the DeepCore region, rather than the full detector. The CNNs use all eight DeepCore strings and the 19 surrounding IceCube strings, which form the two rings outside the DeepCore region, as shown in the orange hexagon in Figure 1. These two rings help optimize reconstruction performance in the DeepCore region by providing additional information from adjacent IceCube DOMs, while avoiding the computational cost of including farther DOMs, which typically do not provide meaningful data for reconstructing neutrino interactions within DeepCore. Due to differences in spatial distance and quantum efficiency between DOMs on the DeepCore and IceCube strings, data from these two categories are fed into separate input layers, each with a dimension of [Nstring, 60, 5], where Nstring represents the number of strings from which data is included (19 and 8 for the IceCube and DeepCore input layers respectively) and 60 corresponds to the 60 DOMs on each string. When the datasets are processed, the input is treated as an image of shape (Nstring × 60) with 5 channels representing the summarized variables (see below) per DOM. The 1D CNN kernel slides along the vertical, or $z$-depth, of the strings with the kernel size given in Figure 2, ranging from taking a maximum of 2 DOMs above and below the target DOM in the convolution. Since the DeepCore strings are deployed in an irregular array, no additional convolution is applied in the $xy$-plane. This is a simplified architecture based on [9], which performs full convolutions over the hexagonally distributed IceCube strings. As an initial attempt to apply CNNs for reconstructing low-energy events in DeepCore, this work demonstrates competitive performance. More recent applications using Graph Neural Networks [6] offer greater generality and are adopted in future analyses. The same convolutional kernel is applied to the IceCube strings to be consistent, and thus both branches of the CNN use the same kernel applied in $z$-depth only. Technically, the 1-D convolution is realized by a “Conv 2D" layer.
After passing through functional layers, such as “Conv 2D", “Batch Normalization", “Max Pooling", and “Dropout" in the sub-networks (see Figure 2 for details), the outputs of the two sub-networks are concatenated and fed into one fully connected dense layer. This last layer is then used to estimate the physical properties, which serve as the labels for the CNN.
@@
The data from each DOM is summarized from the charge and time information of the digitized pulses recorded in the time window of the event interaction, typically within -500 to 4000 ns of the DeepCore trigger time [7]. This time window corresponds to approximately eight times the scattering length divided by the speed of light in ice, so pulses falling outside this window are generally considered noise. To ensure consistency across all data and simulations, we quantize the charges in steps of 0.05 PE before using them to calculate the CNN input features. Five summary variables are calculated from the pulse series for each DOM to represent the statistical features of the digitized waveform: total charge, time of the first pulse, time of the last pulse, charge weighted mean of pulse times, and charge weighted standard deviation of pulse times.
These summary variables have significantly different characteristic scales. To train the networks more efficiently, we rescale the input variables by factors, which are listed in Table 1. These are chosen such that the variables remain near the range [-1,1].
@@
In the output layer, the CNNs use linear activation functions for estimating arrival direction (zenith angle) and interaction vertex coordinates ($x$, $y$, $z$), and rectified linear units (ReLU) for predicting neutrino energy. The difference is because zenith angle, in units of radians, and vertex coordinates, in units of meters, do not have hard boundaries while energy must be positive. For the binary classifications, a binary entropy loss function was used. The two classifiers have different classification tasks and therefore employ different training samples, see Sec. 4 for details, and cannot be combined into a multi-classifier.
@@ 3 Loss function for CNN training
Table 2: Loss functions used for each of the networks.
@@
where $n$ is the number of events per batch, $t_{i}$ is the true value and $r_{i}$ is the reconstructed (or estimated) variable by the CNN. This was chosen carefully such that the energy network would give equal importance to reconstructing the low-energy events (at a few GeV) as the high-energy events (at 100 GeV). Since the $\mathcal{O}$(10) GeV events are the region where oscillation is expected, the percentage error ensures that the 10 GeV events are well-reconstructed without over-emphasis on the $\mathcal{O}$(100) GeV events. The vertex and zenith networks use the mean squared error loss:
@@
The CNNs for classification purposes use a binary cross-entropy loss function:
$\textrm{Loss}=-\frac{1}{n}\sum_{i=1}^{n}{(t_{i}\ln(r_{i})+(1-t_{i})\ln(1-r_{i}))},$ (3.3)
where $t_{i}$ is the MC truth label for event $i$ (1 for signal, 0 for background), and $r_{i}$ is the predicted probability of being a signal event.
@@ 4 Training Samples
Simulated Monte-Carlo (MC) datasets of neutrino interactions and atmospheric muons were used to train the CNNs. In the pre-processing, data quality cuts and random shuffling were applied to ensure the desired distributions were used to train the different CNNs, as discussed further below.
The neutrino MC datasets used in this study for training were generated with the GENIE package (version 2.12.8) [8]. This dataset has larger statistics and is independent of the MC events that were used in the neutrino oscillation data analyses [3]. The final level analysis cuts based on the outcomes of this study are relatively tight compared to the ones that are used as the quality cuts for the training samples. The muon MC samples were generated with MuonGun, an IceCube software package that simulates atmospheric muons from cosmic rays (method based on [11]). The simulated muon events used for training were from a subset of what is used in the oscillation analysis. IceCube neutrino analyses typically involve filters, applied successively, to remove background-like events and keep high-quality neutrino candidates. The training sets for the CNNs were processed through the same sets of filters and basic reconstruction routines commonly used for DeepCore oscillation data analysis (see Section III in [4] for details). However, a boosted decision tree (BDT) [14] was trained using lower-level variables and applied to remove neutrino-like background events, mostly atmospheric muons, therefore noted as BDTμ (see Section III.C of  [4] for details). To drop the noise-like events and keep the track-like $\nu_{\mu}$ charged-current (CC) events, the score of the BDTμ was required to be at least 0.95 in all of the CNN training datasets, except for energy, zenith, and the atmospheric muon dataset used for training the atmospheric muon classification. To retain enough statistics, the atmospheric muon events used for training were taken from the nominal dataset before the BDTμ filter was applied.
@@
Energy $\nu_{\mu}$ $C$ 9 million 1-500
Zenith $\nu_{\mu}$ $C$ 5 million 1-300
Vertex $\nu_{\mu}$ $C$ 5 million 1-500
@@ 4.2 Neutrino Zenith Angle
[…]
@@ 5 Performance
The performance is evaluated on the MC dataset used in the oscillation analysis described in [3] and listed in Table II in [4], with atmospheric flux, neutrino interactions, and oscillation weights applied. The previous IceCube neutrino oscillation result [4] used the SANTA tool for directional reconstruction and the LEERA tool for energy reconstruction, both described in IV.A of [1]. We compared the performance of our CNN routines with these tools using a common subset of events that passed the selections of both the current and previous oscillation analyses. To show the performance on neutrino energy and zenith angle (see Figures 3 and  4), the cuts on energy or zenith angle are not applied. The SANTA algorithm requires events to pass the selection criteria described in Section 2 of [13], it is only applicable to a small set of high-quality events in the DeepCore detector. A more inclusive, likelihood-based algorithm called RETRO has been developed to provide more comprehensive information, allowing it to be applied to a larger set of events and yielding better performance [13]. However, the performance improvements achieved with RETRO come at the cost of significantly slower processing times compared to SANTA/LEERA. The CNNs presented in this paper, which are designed and optimized for the DeepCore region, address the slow processing time issue but offer comparable performance. A comparison of the runtime for SANTA, LEERA, RETRO and the CNNs discussed in this paper, is given in Table 5. A detailed comparison between CNN and RETRO reconstructions can be found in Appendix A.
@@ 5.1 Comparison to Reconstruction used in the Prior DeepCore Oscillation Result
Besides the improvement in reconstruction performance in the selected analysis samples, the CNNs show an advantage in processing time since they can run on graphics processing units (GPU) (see Table 5). The average time per event was measured by running the reconstruction algorithms on a large sample of MC events. For SANTA and RETRO, the runtime depends on the true neutrino energy; therefore, we report the overall mean reconstruction time averaged over the full true neutrino energy spectrum. This is then compared to the average runtime of running the 5 CNNs on a similarly large sample of events within the same energy range. The GPU used for evaluation was a Tesla V100S with a central processing unit (CPU) of Intel(R) Xeon(R) Platinum 8260 CPU @ 2.40 GHz. The same CPU was used to evaluate the inference speed of the 5 CNNs. The precision of the CNN parameters is float32.
Table 5: Runtime of applying the different reconstructions to the MC sample, where the last two columns are scaled from the first one to get a sensible estimate on a large statistics of production. The evaluation of the 5 CNNs is run sequentially.
@@
[…]
@@ 6 Summary
Ongoing efforts include adapting low-energy CNNs for reconstructing low-energy events, with a focus on azimuth and estimating reconstruction uncertainty. These developments will be valuable for future IceCube analyses in both particle physics and astrophysics. The IceCube collaboration continues to prioritize improving reconstruction resolution and runtime through the development of different algorithms. At TeV energies, a CNN takes the whole detector for cascade reconstruction into account [5]. Focusing on sub-100 GeV energies, we developed a CNN that uses DeepCore and surrounding IceCube strings, indicated by the orange hexagon in Figure 1. Additionally, efforts are underway to develop an atmospheric neutrino sample using graph neural networks [6].
@@ Appendix A Comparison to RETRO Reconstruction
Figure 9 shows the comparisons of CNN and RETRO on a subset of overlapping events that pass all the cuts until the analysis-level, and the likelihood fitting quality cut of RETRO with flux, neutrino interaction cross-section, and oscillation weights applied.
@@
The CNN reconstructs the interaction position in ($x$, $y$, $z$) coordinates, but the analysis cut is applied on depth ($z$) and horizontal radius relative to the position of the IceCube string-36 ($\rho_{36}$). Therefore, we show the performance of $z$, $\rho_{36}$, $\cos(\theta_{\rm{zenith}})$, and energy along the true neutrino energy as in Figure 9. In the $\nu_{\mu}$ disappearance analysis [3], the signal events are $\nu_{\mu}$ CC. To optimize reconstruction performance for these signal events, the CNNs for neutrino energy, zenith angle, and interaction vertex were trained and tested on $\nu_{\mu}$ CC events. Similar evaluation plots for the CNN reconstruction of cascade-like events, using $\nu_{e}$ CC as an example, can be found in Figure 10, which also shows reasonable performance. Among all the performance plots, the asymmetric tails in the vertex $z$ distribution arise because good-quality events, which impact the CNN training more effectively, typically have vertices inside the DeepCore region. In contrast, higher-energy events are more likely to originate outside the lower part of DeepCore, leading the CNN to favor deeper $z$ predictions.